A father racing his son has 1/4 the kinetic energy of the son, who has 1/3 the mass of the father. The father speeds up by 1.6 m/s and then has the same kinetic energy as the son. What are the original speeds of (a) the father and (b) the son?
1
Expert's answer
2019-06-19T09:29:11-0400

Kinetic energy formula:


"E=1\/2*mU^2"

let's mark the speed of father UF, the speed of son - US.

EF=mF*U2F/2

ES=mS*U2S/2

EF=1/4*ES

mS=1/3mF

ES=mF*(UF+1.6)2/2


System of equations:

mS*U2S/2=mF(UF+1.6)2/2 (1)

mF=3*mS (2)

mF*U2F/2=1/4*(mS*U2S)/2 (3)

(2) --> (1)

(2) --> (3)


(1) : mS*U2S/2=3*mS*(UF+1.6)2/2 --> U2S=3*(UF+1.6)2

(3) : 3*mS*U2F/2=1/4*mS*U2S/2 --> 3U2F=U2S/2 --> U2S=6*U2S


6*U2F=3*(UF+1.6)2

2*U2F=U2F+3.2*UF+1.62

UF=1/2*(3.2+/-4.5)

UF=5.45 m/s

US=UF*61/2=13.35 m/s


Answer: UF=5.45 m/s, US=13.35 m/s